# Optimal Stopping with a Predicted Prior

### v1

@@ 2 Preliminaries
Let $\mathbb{N}$ denote the set of positive integers. For any $n\in\mathbb{N}$, define $n]\coloneqq\{1,2,\ldots,n\}$. Let $\mathcal{D}$ be the set of all probability distributions supported on $0,\infty)$. We adopt the standard abbreviations cdf, pdf, and pmf for cumulative distribution function, probability density function, and probability mass function. We often use the same symbol for a distribution and its cdf (e.g., $F$), when clear from context. Unless stated otherwise, expectations and probabilities are taken over the randomness of the values and any internal randomness of the algorithm.
@@ 2.1 Optimal Stopping
We consider the maximization problem, where larger values are preferable. The literature has studied two objectives. Let $x_{\textsc{Opt}}$ denote the maximum value among all $n$ values, i.e., $x_{\textsc{Opt}}\coloneqq\max_{i\in[n]}x_{i}$. The max expectation (MaxExp) objective considers the expected value of the algorithm’s accepted decision, $\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]$. The corresponding competitive ratio is
@@
This coincides with the competitive ratio since the offline optimal choice is always maximum. An algorithm is $\Gamma$-competitive if its competitive ratio is at least $\Gamma$.
@@ 2.2 Optimal Stopping with a Predicted Prior
We assume that the number of values $n$ is known. Depending on what the algorithm knows about the prior distribution $F$, we have different optimal stopping problems. If nothing is known, we have essentially the secretary problem [dynkinOptimumChoiceInstant1963, fergusonWhoSolvedSecretary1989]. [^4] The classical secretary problem considers random order without prior. Nonetheless, correaProphetInequalitiesIID2019 showed that the optimal competitive ratio is still $\frac{1}{\mathrm{e}}$ for an unknown prior. With complete information of $F$, we have the i.i.d. prophet inequality [correaUnknownIIDProphets2021].
@@
Following the literature of algorithms with predictions [mitzenmacher2022algorithms], we consider the consistency and robustness of stopping algorithms.
@@ 2.3 Simplifying Assumptions for Algorithm Analyses
For our algorithmic results, we adopt two standard assumptions from the literature that simplify the analysis and are without loss of generality.
Continuous-time Random-order Model.
Besides the above classical model, which we will refer to as the discrete-time model, we may artificially introduce a random order of arrivals by letting each value $x_{i}$ with $i\in[n]$ arrive in a continuous time horizon $[0,1]$, drawing its arrival time $t_{i}$ independently and uniformly from $[0,1]$. Since the values are drawn i.i.d. from $F$, the problem remains the same.
An algorithm for the discrete-time model also works in the continuous-time model by ignoring the arrival time. The converse is also true. We can simulate the continuous-time model by first drawing $n$ arrival times independently and uniformly from $[0,1]$, and then mapping these arrival times in ascending order to the values in their arrival order.
Full-support Continuous Priors.
For the analysis, we further assume that both $F$ and $\tilde{F}$ are continuous with full support on $0,\infty)$. This assumption is without loss of generality, e.g., by independently perturbing each value with non-negative noise drawn from an exponential distribution with parameter $\varepsilon\to 0$; see correaProphetSecretaryBlind2021 for a similar treatment.
For any full-support continuous distribution $F\in\mathcal{D}$, define its right‑continuous inverse as:
@@ 3 Bi-criteria Algorithms
A value $x$ arriving at time $t$ is best-so-far if $x$ is greater than or equal to the prefix maximum for time $t$.
@@
Bi-criteria Algorithm $\mathcal{A}(\tilde{F},\theta)$ For each value $x_{i}$ arriving at time $t_{i}$ when no value has been accepted, accept $x_{i}$ if: 1. Value $x_{i}$ is best-so-far; and 2. The cdf value of $x_{i}$ w.r.t. $\tilde{F}$ is greater than the threshold at time $t_{i}$, i.e., if $\tilde{F}(x_{i})>\theta(t_{i})$.
@@
Specifically, if $\theta(t)=1$, the algorithm accepts values with probability 0 at time $t$, regardless of how large the values are. If $\theta(t)=0$, the algorithm accepts any value that is best-so-far at time $t$, regardless of the value’s relative standing w.r.t. the predicted prior.
@@
The algorithm $\mathcal{A}$ rejects all values before time $t$ if and only if the prefix maximum $y$ for time $t$ and the arrival time $s$ of $y$ satisfy $\tilde{F}(y)\leq\theta(s)$.
@@
If $\tilde{F}(y)\leq\theta(s)$, then all values $x$ arriving before time $s$ are rejected because the cdf value is below the threshold, as $\tilde{F}(x)\leq\tilde{F}(y)\leq\theta(s)\leq\theta(t)$ for any $t\leq s$. Values arriving between $s$ and $t$ are also rejected because they do not exceed $y$ and therefore cannot be best-so-far.
If $\tilde{F}(y)>\theta(s)$, then $y$ is best-so-far and has cdf value above the threshold. Hence, the algorithm accepts either $y$ or a value before $y$. ∎
@@ 3.1 Examples
Example 1 (Dynkin’s algorithm [dynkinOptimumChoiceInstant1963], continuous-time variant).
Dynkin’s algorithm accepts the first best-so-far value after rejecting the first $\lfloor\frac{n}{\mathrm{e}}\rfloor$ values. The continuous-time variant (e.g., feldmanImprovedCompetitiveRatios2011) accepts the first best-so-far value after time $\frac{1}{\mathrm{e}}$, which corresponds to the threshold function $\theta(t)=1$ for $t\in[0,\frac{1}{\mathrm{e}}]$, and $\theta(t)=0$ for $t\in(\frac{1}{\mathrm{e}},1]$.
Dynkin’s algorithm is $\frac{1}{\mathrm{e}}$-competitive for both MaxProb and MaxExp, without any information about the prior [dynkinOptimumChoiceInstant1963, feldmanImprovedCompetitiveRatios2011]. Hence, it is $\frac{1}{\mathrm{e}}$-consistent and $\frac{1}{\mathrm{e}}$-robust.
Example 2 (Gilbert-Mosteller algorithm [gilbertRecognizingMaximumSequence], continuous-time variant).
Gilbert and Mosteller calculated the optimal thresholds for MaxProb through backward induction with a correct prior, i.e., $\tilde{F}=F$. Their algorithm achieves a competitive ratio of approximately $0.580$, with the following analytical form due to samuels1982: [^5] The competitive ratio presented in samuels1982 is $\mathrm{e}^{-\gamma}+(\mathrm{e}^{\gamma}-\gamma-1)\int_{1}^{\infty}\frac{\mathrm{e}^{-\gamma x}}{x}\,\mathrm{d}x$ (fixing a typographical negative sign before the second term), which equals our expression through transformations.
@@
The continuous-time variant under our framework corresponds to the threshold function given by the following equation:
$\int_{t}^{1}\frac{(1-s+s\theta(t))^{n-1}-\theta(t)^{n-1}}{1-s}\,\mathrm{d}s=\theta(t)^{n-1}~.$
It achieves the same consistency $0.580$ for MaxProb as Gilbert-Mosteller algorithm (as a special case of Theorem 2). However, the algorithm is merely 0-robust. To illustrate, consider the case where the prior $F$ is uniform over $[0,1]$, whereas the predicted prior $\tilde{F}$ is uniform over $[2,3]$. In this case, the algorithm rejects all values, as they have cdf value 0 w.r.t. the predicted prior.
Example 3 (Single-threshold algorithm [ehsaniProphetSecretaryCombinatorial2018], i.i.d. case).
Let $\theta(t)=1-\tfrac{1}{n}$ for $t\in[0,1]$. [^6] The original algorithm sets a constant threshold $\theta$ such that $\operatorname{\mathbf{Pr}}[x_{\textsc{Opt}}\geq\theta]=\frac{1}{\mathrm{e}}$, which corresponds to $F(x_{i})\approx 1-\frac{1}{n}$ for each $i$ in the i.i.d. case.
The single-threshold algorithm achieves a competitive ratio of $1-\frac{1}{\mathrm{e}}\approx 0.632$ for the prophet secretary problem in terms of MaxExp. Hence, it is $(1-\frac{1}{\mathrm{e}})$-consistent for MaxExp. However, it suffers from the same counterexample and is only 0-robust like Gilbert-Mosteller algorithm.
@@ 3.2 Robust Threshold Functions
We will further show in Section 5 that the worst-case threshold $F\circ\tilde{F}^{-1}\circ\theta(t)$ is either always 0 or always $1$ in $\lambda_{1},\lambda_{2}]$. They correspond to a variant of Dynkin’s algorithm that starts to accept best-so-far values at a suboptimal time $\lambda=\lambda_{1}$ or $\lambda_{2}$, rather than the optimal choice $\frac{1}{\mathrm{e}}$. The competitive ratio $\lambda\ln\lambda$ of this algorithm (e.g., feldmanImprovedCompetitiveRatios2011) motivates the following definition of robust threshold functions.
@@
Consider any feasible robustness parameter $0\leq\beta\leq\frac{1}{\mathrm{e}}$. Let $0\leq\lambda_{1}\leq\frac{1}{\mathrm{e}}\leq\lambda_{2}\leq 1$ be the solutions to:
@@
For any threshold function $\theta$, define a corresponding robust threshold function as:
$\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}(t)\coloneqq\begin{cases}1&0\leq t\leq\lambda_{1}~;\\ \theta(t)&\lambda_{1}<t<\lambda_{2}~;\\ 0&\lambda_{2}\leq t\leq 1~.\end{cases}$
@@
For example, Figure 2 shows the robust counterpart of the threshold function from Gilbert-Mosteller algorithm.
@@
Figure 2: Illustration of the robust threshold function corresponding to Gilbert–Mosteller algorithm, with $n=10$, $\beta=\nicefrac{{1}}{{3}}$, $\lambda_{1}=\mathrm{e}^{\mathrm{e}^{\mathcal{W}_{-1}(-\beta)}(-\nicefrac{{1}}{{3}})}\approx 0.220$ and $\lambda_{2}=\mathrm{e}^{\mathrm{e}^{\mathcal{W}_{0}(-\beta)}(-\nicefrac{{1}}{{3}})}\approx 0.538$.

### v2

@@ 2 Preliminaries
Let $\mathbb{N}$ denote the set of positive integers. For any $n\in\mathbb{N}$, define $n]\coloneqq\{1,2,\ldots,n\}$. Let $\mathcal{D}$ be the set of all probability distributions supported on $0,\infty)$. We adopt the standard abbreviations cdf, pdf, and pmf for the cumulative distribution function, probability density function, and probability mass function. We use the same symbol for a distribution and its cdf (e.g., $F$). Unless stated otherwise, expectations and probabilities are taken over the randomness of the values and any internal randomness of the algorithm.
@@ 2.1 Optimal Stopping
We consider the maximization problem, where larger values are preferable. The literature has studied two objectives. Let $x_{\textsc{Opt}}$ denote the maximum value among all $n$ values, i.e., $x_{\textsc{Opt}}\coloneqq\max_{i\in[n]}x_{i}$. The max expectation (MaxExp) objective considers the expectation of the algorithm’s accepted value, $\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]$. The corresponding competitive ratio is
@@
This coincides with the competitive ratio since the offline optimal choice is always the maximum. An algorithm is $\Gamma$-competitive if its competitive ratio is at least $\Gamma$.
@@ 2.2 Optimal Stopping with a Predicted Prior
We assume that the number of values $n$ is known. Depending on what the algorithm knows about the prior distribution $F$, we have different optimal stopping problems. If nothing is known, we have essentially the secretary problem [21, 29]. [^5] The classical secretary problem considers random order without prior. Nonetheless, Correa et al. [14] showed that the optimal competitive ratio is still $\nicefrac{{1}}{{\mathrm{e}}}$ for an unknown prior. With complete information of $F$, we have the i.i.d. prophet inequality [13].
@@
Following the literature of algorithms with predictions [46], we define the consistency and robustness of stopping algorithms.
@@ 2.3 Simplifying Assumptions for Algorithm Analyses
For our algorithmic results, we adopt three standard assumptions from the literature that simplify the analysis and are without loss of generality.
Continuous-time Model.
Besides the above classical model, which we will refer to as the discrete-time model, we also consider an equivalent continuous-time model. In this variant, each value $x_{i}$ is further assigned an arrival time $t_{i}$ drawn independently and uniformly from $[0,1]$, at which time the algorithm observes the value. The induced order of arrivals is a uniformly random permutation of the $n$ values, so the discrete-time and continuous-time formulations are equivalent.
Random Tie-breaking.
We assume the predicted prior $\tilde{F}$ is continuous. This assumption is without loss of generality for our algorithm, since we can apply the random tie-breaking rule by Correa et al. [17].
Full-support Continuous Prior.
For the analysis, we further assume that the prior $F$ is continuous with full support on $0,\infty)$. This assumption is without loss of generality. To see this, we construct a perturbed distribution
[…]
Then $F_{\varepsilon}$ is continuous with full support. Moreover, with probability at most $1-(1-\varepsilon)^{n}\leq n\varepsilon$, the instance observed by the algorithm will be changed. Thus, for any finite $n\in\mathbb{N}$, taking $n\varepsilon\to 0$, the performance is preserved.
For any distribution $F\in\mathcal{D}$, define its right‑continuous inverse as:
@@ 3 Bi-criteria Algorithms
A value $x$ arriving at time $t$ is called a record if $x$ is greater than or equal to the prefix maximum for time $t$.
@@
Bi-criteria Algorithm $\mathcal{A}(\tilde{F},\theta)$ For each value $x_{i}$ arriving at time $t_{i}$ when no value has been accepted, accept $x_{i}$ if: 1. Value $x_{i}$ is a record; and 2. The cdf value of $x_{i}$ w.r.t. $\tilde{F}$ is greater than the threshold at time $t_{i}$, i.e., if $\tilde{F}(x_{i})>\theta(t_{i})$.
@@
Specifically, if $\theta(t)=1$, the algorithm does not accept any values at time $t$, regardless of how large the values are. If $\theta(t)=0$, the algorithm accepts any record at time $t$, regardless of the value’s relative standing w.r.t. the predicted prior.
@@
The algorithm $\mathcal{A}(\tilde{F},\theta)$ rejects all values before time $t$ if and only if the prefix maximum $y$ for time $t$ and the arrival time $s$ of $y$ satisfy $\tilde{F}(y)\leq\theta(s)$.
@@
If $\tilde{F}(y)\leq\theta(s)$, then all values $x$ arriving before time $s$ are rejected because the cdf value is below the threshold, as $\tilde{F}(x)\leq\tilde{F}(y)\leq\theta(s)\leq\theta(t)$ for any $t\leq s$. Values arriving between $s$ and $t$ are also rejected because they do not exceed $y$ and therefore cannot be a record.
If $\tilde{F}(y)>\theta(s)$, then $y$ is a record and has cdf value above the threshold. Hence, the algorithm accepts either $y$ or a value before $y$. ∎
@@ 3.1 Examples
Example 1 (Dynkin’s algorithm [21], continuous-time variant).
Dynkin’s algorithm accepts the first record after rejecting the first $\lfloor\nicefrac{{n}}{{\mathrm{e}}}\rfloor$ values. The continuous-time variant (e.g., Feldman et al. [27]) accepts the first record after time $\nicefrac{{1}}{{\mathrm{e}}}$, which corresponds to the threshold function $\theta(t)=1$ for $t\in[0,\nicefrac{{1}}{{\mathrm{e}}}]$, and $\theta(t)=0$ for $t\in(\nicefrac{{1}}{{\mathrm{e}}},1]$.
Dynkin’s algorithm is $\nicefrac{{1}}{{\mathrm{e}}}$-competitive for both MaxProb and MaxExp, without using any information about the prior [21, 27]. Hence, it is $\nicefrac{{1}}{{\mathrm{e}}}$-consistent and $\nicefrac{{1}}{{\mathrm{e}}}$-robust.
Example 2 (Gilbert-Mosteller algorithm [33], continuous-time variant).
With a correct prior $\tilde{F}=F$, the optimal algorithm for the MaxProb objective takes the bi-criteria form according to backward induction. In this setting, Gilbert and Mosteller calculated the optimal thresholds in the discrete-time model. Their algorithm achieves a competitive ratio of approximately $0.580$, with the following analytical form due to Samuels [52]: [^6] The competitive ratio presented in Samuels [52] is $\mathrm{e}^{-\gamma}+(\mathrm{e}^{\gamma}-\gamma-1)\int_{1}^{\infty}\frac{\mathrm{e}^{-\gamma x}}{x}\,\mathrm{d}x$ (fixing a typographical negative sign before the second term), which equals our expression through transformations.
@@
Using a similar backward induction argument, we compute in Lemma 8 the optimal threshold $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ under our bi-criteria framework for any $n\in\mathbb{N}$. For each $t$, $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ is the unique solution $q$ of:
$\int_{t}^{1}\frac{(1-s+sq)^{n-1}-q^{n-1}}{1-s}\,\mathrm{d}s=q^{n-1}~.$
This construction achieves the same consistency $0.580$ for MaxProb as Gilbert-Mosteller algorithm, and it is exactly the form used in the proof of Theorem 2. However, the algorithm is merely 0-robust. To illustrate, consider the case where the prior $F$ is uniform over $[0,1]$, whereas the predicted prior $\tilde{F}$ is uniform over $[2,3]$. In this case, the algorithm rejects all values, as they have cdf value 0 w.r.t. the predicted prior.
Example 3 (Single-threshold algorithm [22], i.i.d. case).
Let $\theta(t)=1-\tfrac{1}{n}$ for $t\in[0,1]$. [^7] The original algorithm sets a constant threshold $\theta$ such that $\operatorname{\mathbf{Pr}}[x_{\textsc{Opt}}\geq\theta]=\nicefrac{{1}}{{\mathrm{e}}}$, which corresponds to $F(x_{i})\approx 1-\nicefrac{{1}}{{n}}$ for each $i$ in the i.i.d. case.
The single-threshold algorithm achieves a competitive ratio of $1-\nicefrac{{1}}{{\mathrm{e}}}\approx 0.632$ for the prophet secretary problem in terms of MaxExp. Hence, it is $(1-\nicefrac{{1}}{{\mathrm{e}}})$-consistent for MaxExp. However, it suffers from the same counterexample and is only 0-robust like Gilbert-Mosteller algorithm.
@@ 3.2 Robust Threshold Functions
In Section 5, we will further show that the worst-case threshold $F\circ\tilde{F}^{-1}\circ\theta(t)$ is either always 0 or always $1$ in $\lambda_{1},\lambda_{2}]$. They correspond to a variant of Dynkin’s algorithm that starts to accept records at a suboptimal time $\lambda=\lambda_{1}$ or $\lambda_{2}$, rather than the optimal choice $\nicefrac{{1}}{{\mathrm{e}}}$, whose competitive ratio is $\lambda\ln\lambda$ for both MaxProb and MaxExp (e.g., by Feldman, Naor, and Schwartz [27]).
This yields the following robustness guarantee, whose proof is deferred to Section 5.
Lemma 7.
For any predicted prior $\tilde{F}$, any threshold function $\theta$, denote $\lambda_{1}=\sup\{0\leq t\leq 1:\theta(t)=1\}$ and $\lambda_{2}=\inf\{0\leq t\leq 1:\theta(t)=0\}$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big)$ is
$\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}\text{-robust}$
for both MaxProb and MaxExp.
Motivated by Lemma 7, we now define the robust counterpart of any threshold function.
@@
For any $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, for any threshold function $\theta$, define its corresponding robust threshold function as:
$\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}(t)\coloneqq\begin{cases}1&0\leq t\leq\lambda_{1}~;\\ \theta(t)&\lambda_{1}<t\leq\lambda_{2}~;\\ 0&\lambda_{2}<t\leq 1~.\end{cases}$
The robustness guarantee in Lemma 7 depends only on the pair $(\lambda_{1},\lambda_{2})$, which allows us to parameterize robustness directly. Consider any feasible robustness parameter $\beta\in[0,\nicefrac{{1}}{{e}}]$. Let $0\leq\lambda_{1}\leq\nicefrac{{1}}{{\mathrm{e}}}\leq\lambda_{2}\leq 1$ be the solutions to:
@@
Lemma 7 above immediately implies that any algorithm with the form $\mathcal{A}(\tilde{F},\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ is $\beta$-robust for both MaxProb and MaxExp.
@@
As an illustration, Figure 2 shows the robust counterpart of the threshold function from the Gilbert-Mosteller algorithm with robustness $\beta=\nicefrac{{1}}{{3}}$.
@@
Figure 2: Illustration of the robust threshold function corresponding to the Gilbert–Mosteller algorithm, with $n=10$, $\beta=\nicefrac{{1}}{{3}}$, $\lambda_{1}=\mathrm{e}^{\mathcal{W}_{-1}(-\nicefrac{{1}}{{3}})}\approx 0.220$ and $\lambda_{2}=\mathrm{e}^{\mathcal{W}_{0}(-\nicefrac{{1}}{{3}})}\approx 0.538$.
3.3 Optimal Robust Threshold Functions
By Lemma 7, once $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$ is fixed, every robust threshold of the form $\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ is guaranteed to be $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$-robust for both MaxProb and MaxExp. Therefore, under this certified robustness lower bound, we can optimize consistency by freely choosing the threshold values on $(\lambda_{1},\lambda_{2})$. Our main result in this subsection is that, for any $n$, there exist objective-specific optimal thresholds $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ (MaxExp) and $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ (MaxProb), whose robust counterparts $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ and $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ maximize consistency for every $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$.
Lemma 8 (Uniqueness and simultaneous optimality of threshold functions).
For any $n$, any prior distribution $F$ and any objective, either MaxExp or MaxProb, there exists a unique non-increasing threshold function, denoted by $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ for MaxExp and by $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ for MaxProb, such that, for any $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, the corresponding bi-criteria algorithm induced by the robust threshold $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ (for MaxExp) or $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ (for MaxProb) achieves the maximum consistency among all algorithms $\mathcal{A}(F,\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$, where $\theta$ ranges over all threshold functions.
Here, the function $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ is precisely the threshold function used in Example 2.
We now sketch the intuition behind why the bi-criteria algorithm admits simultaneously optimal threshold functions. The proof of Lemma 8 relies on the backward-induction equations and on the existence, uniqueness, and monotonicity of the corresponding optimal thresholds for MaxExp and MaxProb; these technical ingredients are developed in Appendix A.
Proof Sketch.
For a fixed objective and prior $F$, consider the dynamic programming formulation of the bi-criteria algorithm with a (not yet specified) threshold function $\theta$. At each time $t$, the algorithm compares the immediate payoff from accepting the current value $x$ with the continuation payoff from rejecting and following the same threshold rule in the future. The optimal threshold at time $t$ is characterized by an indifference condition between these two options.
More precisely, suppose a value $x$ arrives at time $t$ and is a record. If $x$ is right at the threshold, i.e., $F(x)=\theta(t)$, then the algorithm rejects $x$, and we have the following.
  • Prior-free criterion. All future values smaller than $x$ will be rejected, because they cannot be a record.
  • Prior-dependent criterion. Since $\theta$ is non-increasing in $t$ and $\theta(t)=F(x)$, the first future value larger than $x$ will both exceed the threshold and be a new record.
In sum, the algorithm will accept the first future value that exceeds $x$. The continuation payoff is therefore the expected value (for MaxExp) or success probability (for MaxProb) associated with that future acceptance, which depends only on $t$, $x$, and the prior $F$, and does not depend on the robustness parameters $(\lambda_{1},\lambda_{2})$.
In Appendix A, we derive that for each objective and each time $0\leq t\leq 1$, the indifference (Bellman) equation that characterizes the optimal threshold $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)$ for MaxExp and $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ for MaxProb. There we show that each equation has a unique solution for every $t$, and that the resulting threshold function is non-increasing in $t$. The detailed arguments appear in Lemmas 21 and 24 in the appendix.
Since the continuation payoffs are independent of $(\lambda_{1},\lambda_{2})$ when the predicted prior is correct, the same objective-specific threshold solves the indifference equation of time $t$ for every choice of $\lambda_{1}\leq t\leq\lambda_{2}$, and for any $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, the robust threshold $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ (for MaxExp) or $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ (for MaxProb) maximizes consistency among all $\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ in the bi-criteria framework.
This concludes our proof. ∎
For the MaxProb objective, setting the robustness requirement to $\beta=0$ makes the predicted-prior problem coincide with the classical i.i.d. prophet inequality with the MaxProb objective. By Lemma 8, the threshold $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ is the unique consistency-maximizing threshold for this objective within our continuous-time framework. As promised in Example 2, we show that it achieves the optimal $0.580$ competitive ratio in Section 7, where we prove Theorem 2.
In contrast, for MaxExp, our optimal threshold $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ is not directly comparable to those of optimal i.i.d. prophet inequality algorithms, which rely solely on a threshold-based acceptance criterion without any record-based condition. It remains open whether one can attain the $0.745$ worst-case consistency guarantee within this structural class, or whether any such record-based threshold policy must fall strictly below this benchmark.